Grade 9Mechanics

An object floats in water such that 1/5 of its volume is above the surface of water. When the same object is made to float in impure alcohol, the volume that remains outside the water is 1/20 of total volume. The relative density of sample of alcohol is?

Atharv Sagar Suryawanshi
8 Years agoGrade 9

1 Answer

Arun
8 Years ago
If the fraction in water=alfa and the fraction in alcohol=beta
then RD of alcohol=1-alfa/1-beta=(1-1/5)/(1-1/20)=
= (4/5)/(19/20)
= 16/19